# Democratizing Uncertainty Quantification

### v2

@@ 4.5.4 Optimization benchmarks
Optimization Benchmarks
L2-Sea OPT
Constraint deterministic global optimization problem for the total resistance reduction in calm water at fixed speed.
5 Acknowledgements
J. Bénézech research was supported by the UK Engineering and Physical Sciences Research Council (EPSRC) through the Programme Grant: “Certification of Design: Reshaping the Testing Pyramid” EP/S017038/1. The method is developed and integrated in the Distributed and Unified Numerics Environment (DUNE) https://www.dune-project.org/. This work has made use of the Hamilton HPC Service of Durham University.
R. Scheichl and L. Seelinger gratefully acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG, as well as under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).
Appendix A Appendix: Applications
The following sections present some applications of uncertainty quantification to real-world problems that utilize UM-Bridge and demonstrate at the same time aspects of the work presented in the main part of the paper. Specifically, Section A.1 demonstrates programming language independence by connecting an existing Matlab UQ code to a number of model instances written in Fortran. Section A.2 demonstrates parallel runs of a highly complex numerical solver in a UQ application, greatly benefiting from portable models and separation of concerns between model and UQ experts. Finally, Section A.3 demonstrates a complex UQ application combining fast model approximations running on a workstation and transparently offloaded model runs on a cloud cluster of 2,800 processor cores. Models and UQ problems used in Section A.1 and Section A.3 are available as part of the benchmark library.
A.1 Accelerating a naval engineering application
[…]
A.1.1 Problem description
@@ A.1.3 Computational aspects
A.2 Investigating material defects in composite aero-structures
A.2.1 Problem description
@@ A.2.2 UQ workflow
To solve the UQ problem, we employ the Quasi-Monte Carlo method implemented in QMCPy [109], drawing 1024 samples from QMCPy’s DigitalNetB2 generator. We observed that the main influence on the defect is the position along X (maxima located at the corners of the C section) for both criterion, as shown in Figure 8(d). We consider the material to be capable of initiating failure mechanisms under load when any criterion reaches $1$ (the threshold is represented as a dashed red line in the figure). We observed a different distribution for both criterion with a probability of failure of $20.41\%$ for the tensile criterion and $32.42\%$ for the compressive criterion. This is due to the complex interaction between the global geometric feature and the local material behaviour (the laminated stacking sequence is visible in Figure 3). This interaction is not trivial and cannot be anticipated by looking only at the pristine output fields. This methodology is ideal and necessary to conduct an efficient parametric study of defects in large composite parts.
A.2.3 Computational aspects
@@
A.3 Multilevel Delayed Acceptance for tsunami source inversion
A.3.1 Problem description
We solve a Bayesian inference problem based on a tsunami model, which we make available in the benchmark library (see Section B.2.9). It models the propagation of the 2011 Tohoku tsunami by solving the shallow water equations with wetting and drying. For the numerical solution of the PDE, we apply an ADER-DG method implemented in the ExaHyPE framework [110, 111]. Details on the model and its discretization can be found in [112], and include a model with a smoothed bathymetry data incorporating wetting and drying through the use of a finite volume subcell limiter with polynomial order $2$ and a total of $1.7\cdot 10^{5}$ degrees of freedom ($2187$ spatial); and a model using fully resolved bathymetry data and a limited ADER DG scheme of polynomial degree $2$ with a total of $1.7\cdot 10^{7}$ degrees of freedom ($6561$ spatial). Bathymetry data was obtained from GEBCO https://www.gebco.net/data_and_products/gridded_bathymetry_data/.
A.3.2 UQ workflow
@@ A.3.3 Computational aspects
Appendix B Appendix: Benchmark library
This appendix documents all models and benchmark problems currently part of the UM-Bridge benchmark library, using the three categories defined in Section 4.5.
B.1 Models
@@
B.1.1 Euler-Bernoulli beam
@@ B.1.2 L2-Sea model
[…]
The model returns the total resistance $R$ as well as $4$ geometrical quantities (related to the beam, draft, and sonar dome dimensions) whose value changes when playing with the shape parameters $\mathbf{s}$; we postpone their definition to Section B.4.1.
B.1.3 Tsunami model
This model describes the propagation of the 2011 Tohoku tsunami by solving the shallow water equations, see also Section A.3. For the numerical solution of the PDE, we apply an ADER-DG method implemented in the ExaHyPE framework [129].
@@ B.1.5 Tritium desorption
B.1.6 Agent-based disease transmission model
@@ B.1.7 The membrane model
B.1.8 The cookies problem
Figure 10: Computational domain for the cookies problem, Section B.1.8.
@@ B.1.8 The cookies problem
Table 7: Centers of subdomains for the cookies problem, Section B.1.8.
@@ B.2.1 Analytic functions
Name Type Default
Description
a double 2.0
Transformation parameter
b double 0.2
Transformation parameter
scale double 1.0
Scaling factor applied to the underlying normal distribution’s variance
@@
B.2.2 Membrane inverse benchmark
This benchmark for Bayesian inversion is, like the forward model in Section B.1.7, described in substantial detail in [137]. In short, whereas the forward model maps the $64$ coefficients $\theta\in\mathbb{R}^{64}$ that describe $a(\mathbf{x})=a^{\theta}(\mathbf{x})$ to a vector of measurements of $u(\mathbf{x})$ at $139$ distinct locations, denoted by $z^{\theta}\in\mathbb{R}^{139}$, the inverse problem consists of inferring $\theta$ from a specific set of measurements $\hat{z}$. To this end, we define the probability that a given set of coefficients $\theta$ is correct as
@@ B.2.2 Membrane inverse benchmark
B.2.3 Bayesian calibration of tritium desorption
This benchmark performs Bayesian model calibration of the tritium desorption model outlined in Section B.1.5. The benchmark defines a probability density $\pi:\mathbb{R}^{5}\to\mathbb{R}$ where the inputs $\theta$ are $E_{1},E_{2},E_{3}$ (the detrapping energies used in the model description), and $n_{1},n_{2}$ (the density of intrinsic traps).
@@ B.2.3 Bayesian calibration of tritium desorption
B.2.4 Agent-based disease transmission benchmark
This benchmark uses EMOD [136], a stochastic agent-based disease transmission model, to look at how the reproductive number, $R_{0}$, and the correlation between an individual’s acquisition and transmission correlation can affect the ultimate attack fraction and timing of the outbreak peak. This correlation can greatly affect the size and timing of the outbreak for a fixed reproductive number. The model underlying this benchmark is described in Section B.1.6.
@@ B.2.4 Agent-based disease transmission benchmark
Config
Type
Default
Description
@@
epsilon list
[0.001,
 0.001,
 0.001]
Increment used in scipy optimize approx_fprime to estimate the gradient
@@
B.2.6 Computed tomography benchmark
@@
This benchmark follows the same structure and provides the same features as the 1D Deconvolution benchmark in Section B.2.5, providing a posterior distribution of $\mathbf{x}$ given $\mathbf{b}$. It also employs CUQIpy [^3] https://cuqi-dtu.github.io/CUQIpy/ [145, 146] to specify the linear forward model representing CT and uses the same Gaussian noise model.
@@
B.2.7 One dimensional heat Bayesian inverse problem benchmark
@@
[…]
B.2.8 Beam inference
@@
B.2.9 Tsunami source inference
The aim of this benchmark is to obtain the parameters describing the initial displacements leading to the Tohoku tsunami from the data of two available buoys located near the Japanese coast. It is based on the model described in Section B.1.3, which also describes the different levels available. The likelihood of a set of parameters given the simulation results is computed using a weighted average of the maximal wave height and the time at which it is reached. The likelihood is given by a normal distribution with mean given by maximum waveheight and the time at which it is reached for the the two DART buoys 21418 and 21419 (This data can be obtained from NDBC). The covariance matrix shown in Figure 11 depends on the level, but not the probe point.
@@ B.2.9 Tsunami source inference
B.2.10 Coefficient field inversion in a two-dimensional Poisson linear PDE
This is an implementation of the Bayesian inverse problem presented in [150]. The inverse problem estimates a spatially varying diffusion coefficient in an elliptic PDE given limited noisy observations of the PDE solution – the benchmark is thus related to the one discussed in Section B.2.2, but uses a different parameterization. The coefficient in the equation is discretized with finite elements and efficient derivative information is available through the adjoint techniques implemented in hIPPYlib and hIPPYlib-MUQ.
@@ B.2.10 Coefficient field inversion in a two-dimensional Poisson linear PDE
B.2.11 Coefficient field inversion in a Robin boundary condition for a three-dimensional $p$-Poisson nonlinear PDE
@@
The output of the benchmark is a probability that maps a discretized version of $m(x)$ at 233,389 nodes to the product of a prior and a likelihood. The prior distribution on the boundary condition $m(x)$ is a Gaussian process defined through a stochastic differential equation; see [150, Section 5.1.2] for details. The likelihood is formed from observations of $u(x)$ at 300 points $x^{i}$ drawn uniformly over the top surface of the domain. Gaussian noise with a standard deviation of $0.005$ is assumed. Like the benchmark of the previous section, model, prior, likelihood, and sampling of the posterior, as well as convergence diagnostics are implemented with the hIPPYlib-MUQ package.
@@ B.3 Propagation benchmarks
B.3.1 Euler-Bernoulli beam
This is an uncertainty propagation problem modeling the effect of uncertain material parameters on the displacement of an Euler-Bernoulli beam with a prescribed load. Refer to Section B.2.8 for details.
@@ B.3.1 Euler-Bernoulli beam
Mapping Dimensions
Description
input [3]
The value of the beam stiffness in each lumped region.
output [31]
The resulting beam displacement at 31 equidistant grid nodes.
@@
B.3.2 Genz integration and surrogate benchmark
@@
Quadratic decay $\hat{c}_{i}$ $=\frac{1}{(n+1)^{2}}$
Quartic decay $\hat{c}_{i}$ $=\frac{1}{(n+1)^{4}}$
@@
Squared-exponential decay $\hat{c}_{i}$ $=10^{\left(\log_{10}(c_{\mathrm{min}})\frac{(i+1)^{2}}{n}\right)}$
@@
Config Type Default
Description
$n$ integer 1
Function dimension
$C$ float 1
Normalizing constant
$W$ float 0.5
Shift parameter
$T$ string sqexp
coefficient decay type
$N$ string oscillatory
integrand family
@@ B.3.3 L2-Sea propagation
B.3.4 The cookies problem
This benchmark runs a forward UQ problem for the cookies model described in Section B.1.8. The results have been obtained using the Sparse Grids Matlab Kit [153] and the Matlab interface to UM-Bridge. More specifically, we assume that the uncertain parameters $\theta_{n}$ appearing in the definition of the diffusion coefficient Eq. 7 are uniform i.i.d. random variables on the range $[-0.99,-0.2]$ and we aim at computing the expected value of the quantity of interest $\Psi$ in Eq. 8. The structure of this benchmark therefore identical to the test discussed in [139]; however, raw numbers are different since in that work the mesh and PDE solver employed were different (standard FEM with piecewise linear basis on a triangular mesh); in particular, for this benchmark, we fix the splines degree at $p=4$ and the fidelity config parameter to $2$.

### v4

@@ 4.4 Optimization benchmarks
L2-Sea OPT Constraint deterministic global optimization problem for the total resistance reduction in calm water at fixed speed.
5 Applications
The following sections present some applications of uncertainty quantification to real-world problems that utilize UM-Bridge and demonstrate at the same time aspects of the work presented in the main part of the paper. Specifically, Section 5.1 demonstrates programming language independence by connecting an existing Matlab UQ code to a number of model instances written in Fortran. Section 5.2 demonstrates parallel runs of a highly complex numerical solver in a UQ application, greatly benefiting from portable models and separation of concerns between model and UQ experts. Finally, Section 5.3 demonstrates a complex UQ application combining fast model approximations running on a workstation and transparently offloaded model runs on a cloud cluster of 2,800 processor cores. Models and UQ problems used in Section 5.1 and Section 5.3 are available as part of the benchmark library.
5.1 Accelerating a naval engineering application
5.1.1 Problem description
@@ 5.1.3 Computational aspects
5.2 Investigating material defects in composite aero-structures
5.2.1 Problem description
@@ 5.2.2 UQ workflow
To solve the UQ problem, we employ the Quasi-Monte Carlo method implemented in QMCPy [14], drawing 1024 samples from QMCPy’s DigitalNetB2 generator. We observed that the main influence on the defect is the position along X (maxima located at the corners of the C section) for both criterion, as shown in Figure 7(d). We consider the material to be capable of initiating failure mechanisms under load when any criterion reaches $1$ (the threshold is represented as a dashed red line in the figure). We observed a different distribution for both criterion with a probability of failure of $20.41\%$ for the tensile criterion and $32.42\%$ for the compressive criterion. This is due to the complex interaction between the global geometric feature and the local material behaviour (the laminated stacking sequence is shown on the right-hand side of Figure 2). This interaction is not trivial and cannot be anticipated by looking only at the pristine output fields. This methodology is ideal and necessary to conduct an efficient parametric study of defects in large composite parts.
5.2.3 Computational aspects
@@
5.3 Multilevel Delayed Acceptance for tsunami source inversion
5.3.1 Problem description
We solve a Bayesian inference problem based on a tsunami model, which we make available in the benchmark library (see Section A.2.9). It models the propagation of the 2011 Tohoku tsunami by solving the shallow water equations with wetting and drying. For the numerical solution of the PDE, we apply an ADER-DG method implemented in the ExaHyPE framework [45, 46]. Details on the model and its discretization can be found in [47], and include a model with a smoothed bathymetry data incorporating wetting and drying through the use of a finite volume subcell limiter with polynomial order $2$ and a total of $1.7\cdot 10^{5}$ degrees of freedom ($2187$ spatial); and a model using fully resolved bathymetry data and a limited ADER DG scheme of polynomial degree $2$ with a total of $1.7\cdot 10^{7}$ degrees of freedom ($6561$ spatial). Bathymetry data was obtained from GEBCO https://www.gebco.net/data_and_products/gridded_bathymetry_data/.
5.3.2 UQ workflow
@@ 5.3.3 Computational aspects
6 Discussion
Package Link Scalability
Multiplicity Model language agnostic Technology
hIPPYlib-MUQ [52] one-to-one No C++, Python Single node
MUQ [12] one-to-many No C++, Python HPC (MPI)
PyMC [7], QMCPy [14] one-to-many No Python Single node
Dakota [53] one-to-many Yes C++, Files HPC (SLURM)
lagun [54] one-to-many Yes Files HPC (SLURM)
UQLab [55] one-to-many Yes Matlab, files HPC (SLURM)
UM-Bridge [25] many-to-many Yes C++, Julia, Matlab, Python, R HPC (SLURM, PBS), Kubernetes
Limitations
7 Outlook
8 Acknowledgements
J. Bénézech research was supported by the UK Engineering and Physical Sciences Research Council (EPSRC) through the Programme Grant EP/S017038/1 “Certification of Design: Reshaping the Testing Pyramid” and made use of the Hamilton HPC Service of Durham University. The code is developed and integrated in the Distributed and Unified Numerics Environment (DUNE) https://www.dune-project.org/.
R. Scheichl and L. Seelinger gratefully acknowledge support by the state of Baden-Württemberg through bwHPC, as well as the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG, as well as its Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).
Appendix A Appendix: Benchmark library
This appendix documents all models and benchmark problems currently part of the UM-Bridge benchmark library, using the three categories defined in Section 4.
A.1 Models
@@
A.1.1 Euler-Bernoulli beam
@@ A.1.2 L2-Sea model
Moreover, three configuration parameters can be set to control the behavior for this model: fidelity is an integer value that controls the mesh refinement, ranging from to 1 (highest resolution, default) to 7 (lowest resolution), with a refinement ratio of 20.25, see [39] for more details; sinkoff is a character that can be set to ’y’ (default) / ’n’ to respectively disable/enable sinkage secondary movements (up/down translation on the vertical axis); similarly, trimoff can be set to ’y’ (default) / ’n’ to respectively disable/enable trim secondary movement (up/down rotation about its transverse axis, i.e. bow to stern).
The model returns the total resistance $R$ as well as $4$ geometrical quantities (related to the beam, draft, and sonar dome dimensions) whose value changes when playing with the shape parameters $\mathbf{s}$; we postpone their definition to Section A.4.1.
A.1.3 Tsunami model
This model describes the propagation of the 2011 Tohoku tsunami by solving the shallow water equations, see also Section 5.3. For the numerical solution of the PDE, we apply an ADER-DG method implemented in the ExaHyPE framework [73].
@@ A.1.5 Tritium desorption
A.1.6 Agent-based disease transmission model
@@ A.1.7 The membrane model
A.1.8 The cookies problem
Figure 9: Computational domain for the cookies problem, Section A.1.8.
@@ A.1.8 The cookies problem
Table 7: Centers of subdomains for the cookies problem, Section A.1.8.
@@ A.2.1 Analytic functions
Name Type Default Description
a double 2.0 Transformation parameter
b double 0.2 Transformation parameter
scale double 1.0 Scaling factor applied to the underlying normal distribution’s variance
@@
A.2.2 Membrane inverse benchmark
This benchmark for Bayesian inversion is, like the forward model in Section A.1.7, described in substantial detail in [21]. In short, whereas the forward model maps the $64$ coefficients $\theta\in\mathbb{R}^{64}$ that describe $a(\mathbf{x})=a^{\theta}(\mathbf{x})$ to a vector of measurements of $u(\mathbf{x})$ at $139$ distinct locations, denoted by $z^{\theta}\in\mathbb{R}^{139}$, the inverse problem consists of inferring $\theta$ from a specific set of measurements $\hat{z}$. To this end, we define the probability that a given set of coefficients $\theta$ is correct as
@@ A.2.2 Membrane inverse benchmark
A.2.3 Bayesian calibration of tritium desorption
This benchmark performs Bayesian model calibration of the tritium desorption model outlined in Section A.1.5. The benchmark defines a probability density $\pi:\mathbb{R}^{5}\to\mathbb{R}$ where the inputs $\theta$ are $E_{1},E_{2},E_{3}$ (the detrapping energies used in the model description), and $n_{1},n_{2}$ (the density of intrinsic traps).
@@ A.2.3 Bayesian calibration of tritium desorption
A.2.4 Agent-based disease transmission benchmark
This benchmark uses EMOD [80], a stochastic agent-based disease transmission model, to look at how the reproductive number, $R_{0}$, and the correlation between an individual’s acquisition and transmission correlation can affect the ultimate attack fraction and timing of the outbreak peak. This correlation can greatly affect the size and timing of the outbreak for a fixed reproductive number. The model underlying this benchmark is described in Section A.1.6.
@@ A.2.4 Agent-based disease transmission benchmark
Config Type Default Description
@@
epsilon list [0.001,  0.001,  0.001] Increment used in scipy optimize approx_fprime to estimate the gradient
@@
A.2.6 Computed tomography benchmark
@@
This benchmark follows the same structure and provides the same features as the 1D Deconvolution benchmark in Section A.2.5, providing a posterior distribution of $\mathbf{x}$ given $\mathbf{b}$. It also employs CUQIpy [^3] https://cuqi-dtu.github.io/CUQIpy/ [10, 9] to specify the linear forward model representing CT and uses the same Gaussian noise model.
@@
A.2.7 One dimensional heat Bayesian inverse problem benchmark
@@
Figure 15: The small noise case of the 1d inverse heat benchmark. (a) Noisy data, exact data (solution at $\tau=\tau^{\text{max}}$), and exact solution (solution at $\tau=0$). Note that for the small noise case, noisy and exact data are visually indistinguishable. (b) Samples’ $95\%$ credible interval computed after mapping the samples of KL coefficients $\mathbf{x}$ to samples of the corresponding function $\mathbf{g}$. (c) KL coefficients samples’ $95\%$ credible intervals.
A.2.8 Beam inference
@@
A.2.9 Tsunami source inference
The aim of this benchmark is to obtain the parameters describing the initial displacements leading to the Tohoku tsunami from the data of two available buoys located near the Japanese coast. It is based on the model described in Section A.1.3, which also describes the different levels available. The likelihood of a set of parameters given the simulation results is computed using a weighted average of the maximal wave height and the time at which it is reached. The likelihood is given by a normal distribution with mean given by maximum waveheight and the time at which it is reached for the the two DART buoys 21418 and 21419 (This data can be obtained from NDBC). The covariance matrix shown in Table 11 depends on the level, but not the probe point.
@@ A.2.9 Tsunami source inference
A.2.10 Coefficient field inversion in a two-dimensional Poisson linear PDE
This is an implementation of the Bayesian inverse problem presented in [52]. The inverse problem estimates a spatially varying diffusion coefficient in an elliptic PDE given limited noisy observations of the PDE solution – the benchmark is thus related to the one discussed in Section A.2.2, but uses a different parameterization. The coefficient in the equation is discretized with finite elements and efficient derivative information is available through the adjoint techniques implemented in hIPPYlib and hIPPYlib-MUQ.
@@ A.2.10 Coefficient field inversion in a two-dimensional Poisson linear PDE
A.2.11 Coefficient field inversion in a Robin boundary condition for a three-dimensional $p$-Poisson nonlinear PDE
@@
The output of the benchmark is a probability that maps a discretized version of $m(x)$ at 233,389 nodes to the product of a prior and a likelihood. The prior distribution on the boundary condition $m(x)$ is a Gaussian process defined through a stochastic differential equation; see [52, Section 5.1.2] for details. The likelihood is formed from observations of $u(x)$ at 300 points $x^{i}$ drawn uniformly over the top surface of the domain. Gaussian noise with a standard deviation of $0.005$ is assumed. Like the benchmark of the previous section, model, prior, likelihood, and sampling of the posterior, as well as convergence diagnostics are implemented with the hIPPYlib-MUQ package.
@@ A.3 Propagation benchmarks
A.3.1 Euler-Bernoulli beam
This is an uncertainty propagation problem modeling the effect of uncertain material parameters on the displacement of an Euler-Bernoulli beam with a prescribed load. Refer to Section A.2.8 for details.
@@ A.3.1 Euler-Bernoulli beam
Mapping Dimensions Description
input [3] The value of the beam stiffness in each lumped region.
output [31] The resulting beam displacement at 31 equidistant grid nodes.
@@
A.3.2 Genz integration and surrogate benchmark
@@
Quadratic decay $\hat{c}_{i}$ $=\frac{1}{(i+1)^{2}}$
Quartic decay $\hat{c}_{i}$ $=\frac{1}{(i+1)^{4}}$
@@
Squared-exponential decay $\hat{c}_{i}$ $=10^{\left(\log_{10}(c_{\mathrm{min}})\frac{(i+1)^{2}}{n^{2}}\right)}$
@@
Config Type Default Description
$n$ integer 1 Function dimension
$C$ float 1 Normalizing constant
$W$ float 0.5 Shift parameter
$T$ string sqexp coefficient decay type
$N$ string oscillatory integrand family
@@ A.3.3 L2-Sea propagation
A.3.4 The cookies problem
This benchmark runs a forward UQ problem for the cookies model described in Section A.1.8. The results have been obtained using the Sparse Grids Matlab Kit [15] and the Matlab interface to UM-Bridge. More specifically, we assume that the uncertain parameters $\theta_{n}$ appearing in the definition of the diffusion coefficient Eq. 7 are uniform i.i.d. random variables on the range $[-0.99,-0.2]$ and we aim at computing the expected value of the quantity of interest $\Psi$ in Eq. 8. The structure of this benchmark therefore identical to the test discussed in [82]; however, raw numbers are different since in that work the mesh and PDE solver employed were different (standard FEM with piecewise linear basis on a triangular mesh); in particular, for this benchmark, we fix the splines degree at $p=4$ and the fidelity config parameter to $2$.